# OpenEarthAgent[Uncaptioned image]: A Unified Framework for Tool-Augmented Geospatial Agents

### v1

@@ Abstract
Recent progress in multimodal reasoning has enabled agents that can interpret imagery, connect it with language, and perform structured analytical tasks. Extending such capabilities to the remote sensing domain remains challenging, as models must reason over spatial scale, geographic structures, and multispectral indices while maintaining coherent multi-step logic. To bridge this gap, OpenEarthAgent introduces a unified framework for developing tool-augmented geospatial agents trained on satellite imagery, natural-language queries, and detailed reasoning traces. The training pipeline relies on supervised fine-tuning over structured reasoning trajectories, aligning the model with verified multi-step tool interactions across diverse analytical contexts. The accompanying corpus comprises 14,538 training and 1,169 evaluation instances, with more than 100K reasoning steps in the training split and over 7K reasoning steps in the evaluation split. It spans urban, environmental, disaster, and infrastructure domains, and incorporates GIS-based operations alongside index analyses such as NDVI, NBR, and NDBI. Grounded in explicit reasoning traces, the learned agent demonstrates structured reasoning, stable spatial understanding, and interpretable behaviour through tool-driven geospatial interactions across diverse conditions. We report consistent improvements over a strong baseline and competitive performance relative to recent open and closed-source models.
Figure 1: Comparison between baseline Qwen3-4B and OpenEarthAgent-4B on a GIS reasoning query. OpenEarthAgent executes tools sequentially with correct dependencies and interpretable feedback, while Qwen3 fails due to unordered multi-tool calls and inconsistent reasoning.
@@ 1 Introduction
The evolution of visual representation learning has transitioned from static perception to interactive multi-modal reasoning. Early single-shot frameworks such as DINO [3] and MAE [13] laid the foundation for large-scale visual understanding through self-supervised objectives, learning strong image encoders without explicit language-based reasoning or interaction. Inspired by these successes, the remote sensing community extended this paradigm to earth observation, where models such as Prithvi [15], Copernicus-FM [40], Galileo [36], Panopticon [37], TerraFM [6], and AnySat [1] scaled representation learning to global multisensor data. These earth-scale vision models demonstrated remarkable transfer across spatial resolutions and modalities, yet their predictions remained one-shot and perception-centric, focusing on recognition rather than structured reasoning.
In parallel, large multi-modal models have progressed beyond perception. Architectures such as BLIP-2 [18], InstructBLIP [5], LLaVA-OneVision [17], and Kosmos-2 [27] extended language models into the visual domain, enabling grounded image understanding. Building on this trajectory, remote-sensing VLMs emerged to adapt multimodal reasoning to geospatial data. Early efforts such as RemoteCLIP [21], SkySenseGPT [23], and GeoChat [16] introduced large-scale multimodal alignment for remote sensing through paired image–text and instruction-following datasets. Recently, EarthDial [31] advanced this paradigm by extending grounding across optical, SAR, thermal, and temporal modalities, enabling unified multimodal reasoning for classification, captioning, and change analysis. Despite coupling language with perception, current remote sensing VLMs remain largely descriptive and lack explicit structured reasoning.
Subsequent reasoning-centric approaches such as ReAct [47], DeepSeek-R1 [10], and VLM-R1 [30] introduced structured planning and, in some cases, tool-driven reasoning, transforming static perception into organized multi-step task execution. Recent developments such as OpenThinkIMG [33] unified these advances through multi-step vision-tool interaction, where large vision-language models iteratively “think with images” through executable reasoning trajectories. Inspired by these advances, efforts such as ThinkGeo [29] and Earth-Agent [7] began exploring tool-augmented reasoning for earth observation (EO). They demonstrate that large models can plan analytical chains but still face challenges in geospatial grounding, coordinate consistency, and physically verifiable outputs. These limitations highlight the need for geographically grounded agents that integrate perception with explicit and interpretable reasoning.
@@
To assess generalization and reasoning fidelity, we benchmark a broad suite of large and reasoning models, including the GPT family [25], Qwen [46], InternLM [2], and other recent multimodal LLMs. Our dataset is constructed through a unified data-gathering pipeline, where each modality undergoes question synthesis and automated validation to ensure reasoning diversity, spatial grounding, and cross-source consistency. This large-scale, tool-augmented corpus bridges the divide between GIS analysis and remote-sensing perception, providing a unified setting for evaluating structured geospatial reasoning. This work introduces an agentic framework for remote sensing that integrates tool-based reasoning within a unified training and evaluation setup.
@@
  • Structured reasoning-alignment framework: A supervised training setup that aligns multi-step tool interactions to ensure consistent reasoning flow and spatially grounded decision making (Fig. 1).
@@ 2 Related Work
Remote sensing has progressed from task-specific models to large-scale foundation and vision-language models trained on global, multimodal data. Early self-supervised methods [24, 4] established transferable pretraining strategies on Sentinel archives. Subsequent foundation models include Prithvi-v2 [35] with spatiotemporal transformers on HLS, Copernicus-FM [40] with metadata-aware hypernetworks for Sentinel, Panopticon [37] and AnySat [1] with spectral- and resolution-adaptive embeddings, and CROMA [8] fusing radar and optical contrastive learning. Galileo [36] captures global-to-local context, demonstrating scalable pretraining across diverse sensors and geographies. On the multimodal side, GeoChat [16] enabled language grounding to satellite imagery, and EarthDial [31] extended this paradigm to multi-resolution optical, SAR, thermal, and temporal modalities across diverse tasks. Despite their scalability and representational strength, these models remain largely single-pass encoders; they lack structured reasoning, tool orchestration, and verifiable intermediate analysis.
@@
Alongside foundation models, agentic systems have emerged that couple reasoning with external tools and feedback-driven control. Early frameworks such as ReAct [47] and Voyager [39] demonstrated the transition from static inference to autonomous, goal-directed reasoning. Subsequent systems, including WebAgent [42], VisTA [14], and DeepEyes [50], introduced modular architectures capable of selecting and sequencing tools under guided supervision, improving reasoning depth and execution consistency. OpenThinkIMG [33] unified these ideas through large-scale visual–tool integration with standardized APIs, while OctoTools [22] emphasized modular, verifiable execution via structured tool interfaces. Collectively, these systems established key principles: looped reasoning, standardized tool I/O, and trajectory-based learning, yet remain geospatially naive, lacking coordinate awareness, scale handling, and domain-specific spatial verification essential for earth observation analysis.
@@
An emerging wave of EO research is now bringing these agentic principles into geospatial contexts. ThinkGeo [29] frames EO question answering as tool-augmented reasoning, benchmarking ReAct-style chains and exposing persistent weaknesses in coordinate consistency, spatial grounding, and multi-step planning. Earth-Agent [7] broadens the tool ecosystem to include spectral products and standardized interfaces but still relies on largely predefined workflows, limiting physically verifiable GIS and index-based reasoning. Geo-OLM [32] explores plan-tool-verify prompting for compact models, improving efficiency but depending primarily on prompt-level heuristics rather than learned adaptive policies. Building on this gap, we introduce an EO-native agentic stack that unifies dataset construction, supervised reasoning alignment, and evaluation under a physically grounded framework. OpenEarthAgent enables broad orchestration of GIS, index-based, optical, and SAR operations through structured instruction-tool pairs; trains on detailed reasoning traces with a dedicated evaluation set; and supports interpretable, verifiable, multi-step reasoning across diverse geospatial conditions.
3 Dataset Curation
@@
ReAct [47]
OpenThinkIMG [33]
ThinkGeo [29]
Earth-Agent [7]
Geo-OLM [32]
RS-Agent [45]
RS-ChatGPT [11]
@@
DIOR [19] B-Box, Category 0.5–30 2024 Transport & Aviation Monitoring; Infrastructure Mapping
DOTA [43] GSD, B-Box, Category 0.1–1 2021 Transport Surveillance; Infrastructure Assessment
NWPU-VHR-10 [38] B-Box, Category 0.5–2 2023 Aviation; Transport; Infrastructure
UCAS-AOD [51] B-Box, Category 0.5–2 2015 Aviation; Transport
AID [44] B-Box, Category 0.2–2 2017 Urban Planning; Industrial Site Detection
iSAID [41] GSD, B-Box, Seg. Map, Pix. Count 0.1–1 2019 Segmentation; Transport Monitoring
xBD [12] GSD, B-Box, Category, Pix. Count 1–3.5 2019 Disaster Assessment; Change Detection
FloodNet [28] GSD, B-Box, Category, Seg. Map, Pix. Count 0.015–0.02 2020 Flood Monitoring; Damage Estimation
Global-Dumpsite [34] B-Box, Category 0.3–0.8 2023 Environmental Monitoring; Waste Localization
@@
SSDD [49] B-Box, Category 1–15 2021 Ship Detection; Maritime Transport
SADD [48] B-Box, Category 0.5–3 2022 Aviation; Vessel Identification
SIVED [20] B-Box, Category 0.1–0.3 2023 Vehicle Detection; Transport Monitoring
@@
OpenStreetMap [26] POIs (Points of Interest) N/A 2017- Urban Mapping; Infrastructure Analysis
@@
Google Earth Engine [9] Index Layers 10–30 2017- Urban Planning; Change Detection; Environmental Monitoring
@@
Data Sources and Composition: The dataset integrates heterogeneous imagery and metadata from open-access EO repositories, encompassing optical, SAR, GIS, and multispectral domains. Primary sources include high-resolution benchmarks such as DOTA [43], DIOR [19], xBD [12], AID [44], NWPU-VHR-10 [38], FloodNet [28], and Global-Dumpsite [34], complemented by Sentinel collections and Google Earth Engine archives (see Table 2 for details). Each dataset offers distinct spatial resolutions and annotation formats (bounding boxes, segmentation masks, category labels), providing both fine-grained object understanding and large-scale semantic context. The collection spans seven thematic domains: urban planning, disaster assessment, environmental monitoring, transportation, aviation, recreation, and industrial infrastructure. Multispectral and indexed layers (e.g., NDVI, NBR, NDBI) are derived through Google Earth Engine to complement optical and SAR imagery with physically meaningful environmental indicators. Where required for spatial reasoning tasks, samples are associated with geospatial metadata such as coordinates, projection reference, and ground sampling distance (GSD), enabling verifiable geospatial analysis.
@@
Figure 4: Key statistics for the curated corpus used for training and evaluation of our proposed agentic pipeline.
Dataset Statistics: Fig. 4 summarizes the scale and structure of the OpenEarthAgent dataset. The training split contains 14,538 instances and the benchmark split contains 1,169 evaluation samples. Each instance includes multimodal imagery, a natural-language query, and a structured reasoning trace with interleaved tool calls and intermediate observations. Across the training set, the corpus contains 100,656 reasoning steps (average 6.92 per query)and the benchmark split contains 7,064 reasoning steps (average 6.04 per query). Reasoning steps correspond to explicit thought-action-observation transitions, while tool calls capture the diversity of operational pathways required for solving curated agentic tasks. The breadth of modalities, thematic coverage, and depth of reasoning traces enable rigorous evaluation of grounded, interpretable, and tool-augmented geospatial reasoning.
4 Methodology
4.1 Overview
@@
4.2 Unified Tool Registry
@@
4.3 Reasoning Trajectories and Training Objective
@@ 5.1 Implementation Details
OpenEarthAgent is finetuned using Qwen3-4B-Instruct-2507 as the base model. Training was performed on LLM controller on four NVIDIA A100 GPUs (40 GB). The base model is loaded via Unsloth’s FastLanguageModel, enabling efficient multi-GPU distributed training. The model is trained for one epoch with a learning rate of $2\times 10^{-5}$, cosine learning rate scheduling, 0.05 warmup ratio, and a batch size of 16. and a batch size of 16. The maximum sequence length is set to 4096 tokens to support long-context conversational tool interactions.

### v2

@@ Abstract
Recent progress in multimodal reasoning has enabled agents that interpret imagery, connect it with language, and execute structured analytical tasks. Extending these capabilities to remote sensing remains challenging, as models must reason over spatial scale, geographic structures, and multispectral indices while maintaining coherent multi-step logic. To address this gap, we introduce OpenEarthAgent, a unified framework for tool-augmented geospatial reasoning trained on satellite imagery, natural-language queries, and structured reasoning traces. Beyond serving as a benchmark, OpenEarthAgent establishes a cohesive agentic architecture built around a unified executable tool registry and trajectory-based policy learning. The framework standardizes heterogeneous visual, spectral, GIS, and georeferenced raster operations under a consistent callable schema, enabling modular orchestration and deterministic execution. Training is performed via supervised fine-tuning on structured reasoning trajectories, coupled with deterministic replay validation to ensure full-chain executability and spatial correctness. The accompanying corpus comprises 14,538 training and 1,169 evaluation instances with over 107K reasoning steps, spanning urban, environmental, disaster, and infrastructure domains and incorporating GIS-based operations alongside index analyses such as NDVI, NBR, and NDBI. Grounded in explicit multi-step traces, the learned agent exhibits structured reasoning, stable spatial grounding, and interpretable tool-driven behavior across diverse EO scenarios. We report consistent improvements over a strong baseline and competitive performance relative to recent open and closed-source models. Grounded in explicit reasoning traces, the learned agent demonstrates structured reasoning, stable spatial understanding, and interpretable behaviour through tool-driven geospatial interactions across diverse conditions. We report consistent improvements over a strong baseline and competitive performance relative to recent open and closed-source models.
Figure 1: Comparison of OpenEarthAgent-4B and Qwen3-4B on a complex GIS reasoning task. OpenEarthAgent correctly sequences tool calls with proper dependencies and feedback, while Qwen3 fails due to misordered tool usage and inconsistent reasoning.
@@ 1 Introduction
The evolution of visual representation learning has transitioned from static perception to interactive multi-modal reasoning. Early single-shot frameworks such as DINO [caron2021emerging] and MAE [he2022masked] laid the foundation for large-scale visual understanding through self-supervised objectives, learning strong image encoders without explicit language-based reasoning or interaction. Inspired by these successes, the remote sensing community extended this paradigm to earth observation, where models such as Prithvi [jakubik2023foundation], Copernicus-FM [wang2025towards], Galileo [tseng2025galileo], Panopticon [waldmann2025panopticon], TerraFM [danish2025terrafm], and AnySat [astruc2025anysat] scaled representation learning to global multisensor data. These earth-scale vision models demonstrated remarkable transfer across spatial resolutions and modalities, yet their predictions remained one-shot and perception-centric, focusing on recognition rather than structured reasoning.
In parallel, large multi-modal models have progressed beyond perception. Architectures such as BLIP-2 [li2023blip], InstructBLIP [dai2023instructblip], LLaVA-OneVision [li2024llava], and Kosmos-2 [peng2023kosmos] extended language models into the visual domain, enabling grounded image understanding. Building on this trajectory, remote-sensing VLMs emerged to adapt multimodal reasoning to geospatial data. Early efforts such as RemoteCLIP [liu2024remoteclip], SkySenseGPT [luo2024skysensegpt], and GeoChat [kuckreja2024geochat] introduced large-scale multimodal alignment for remote sensing through paired image–text and instruction-following datasets. Recently, EarthDial [soni2025earthdial] advanced this paradigm by extending grounding across optical, SAR, thermal, and temporal modalities, enabling unified multimodal reasoning for classification, captioning, and change analysis. Despite coupling language with perception, current remote sensing VLMs remain largely descriptive and lack explicit structured reasoning.
Subsequent reasoning-centric approaches such as ReAct [yao2022react], DeepSeek-R1 [guo2025deepseek], and VLM-R1 [shen2025vlm] introduced structured planning and, in some cases, tool-driven reasoning, transforming static perception into organized multi-step task execution. Recent developments such as OpenThinkIMG [su2025openthinkimg] unified these advances through multi-step vision-tool interaction, where large vision-language models iteratively “think with images” through executable reasoning trajectories. Inspired by these advances, efforts such as ThinkGeo [shabbir2025thinkgeo] and Earth-Agent [feng2025earth] began exploring tool-augmented reasoning for earth observation (EO). They demonstrate that large models can plan analytical chains but still face challenges in geospatial grounding, coordinate consistency, and physically verifiable outputs. These limitations highlight the need for geographically grounded agents that integrate perception with explicit and interpretable reasoning.
@@
Central to this design is a unified tool registry that abstracts heterogeneous geospatial operators through a standardized executable schema, enabling consistent orchestration across perception, GIS computation, spectral analysis, and georeferenced raster reasoning. Unlike prior EO benchmarks that focus primarily on data scale, our contribution lies in coupling the dataset with an execution engine and trajectory-based policy alignment. This integration ensures that reasoning steps are not only generated but deterministically validated for spatial and logical consistency. To assess generalization and reasoning fidelity, we benchmark a broad suite of large and reasoning models, including the GPT family [OpenAI2025GPT], Qwen [yang2025qwen3], InternLM [cai2024internlm2], and other recent multimodal LLMs. Our dataset is constructed through a unified data-gathering pipeline, where each modality undergoes question synthesis and automated validation to ensure reasoning diversity, spatial grounding, and cross-source consistency. This large-scale, tool-augmented corpus bridges the divide between GIS analysis and remote-sensing perception, providing a unified setting for evaluating structured geospatial reasoning. This work introduces an agentic framework for remote sensing that integrates tool-based reasoning within a unified training and evaluation setup.
@@
  • Agentic reasoning framework with deterministic validation: A unified tool registry and a trajectory-based supervised alignment mechanism that enforces consistent tool syntax, executable reasoning chains, and spatially grounded decision-making (Fig. 1).
@@ 2 Related Work
Remote sensing has progressed from task-specific models to large-scale foundation and vision-language models trained on global, multimodal data. Early self-supervised methods [manas2021seasonal, cong2022satmae] established transferable pretraining strategies on Sentinel archives. Subsequent foundation models include Prithvi-v2 [szwarcman2024prithvi] with spatiotemporal transformers on HLS, Copernicus-FM [wang2025towards] with metadata-aware hypernetworks for Sentinel, Panopticon [waldmann2025panopticon] and AnySat [astruc2025anysat] with spectral- and resolution-adaptive embeddings, and CROMA [fuller2023croma] fusing radar and optical contrastive learning. Galileo [tseng2025galileo] captures global-to-local context, demonstrating scalable pretraining across diverse sensors and geographies. On the multimodal side, GeoChat [kuckreja2024geochat] enabled language grounding to satellite imagery, and EarthDial [soni2025earthdial] extended this paradigm to multi-resolution optical, SAR, thermal, and temporal modalities across diverse tasks. Despite their scalability and representational strength, these models remain largely single-pass encoders; they lack structured reasoning, tool orchestration, and verifiable intermediate analysis.
@@
Alongside foundation models, agentic systems have emerged that couple reasoning with external tools and feedback-driven control. Early frameworks such as ReAct [yao2022react] and Voyager [wang2023voyager] demonstrated the transition from static inference to autonomous, goal-directed reasoning. Subsequent systems, including WebAgent [wei2025webagent], VisTA [huang2025visualtoolagent], and DeepEyes [zheng2025deepeyes], introduced modular architectures capable of selecting and sequencing tools under guided supervision, improving reasoning depth and execution consistency. OpenThinkIMG [su2025openthinkimg] unified these ideas through large-scale visual–tool integration with standardized APIs, while OctoTools [lu2025octotools] emphasized modular, verifiable execution via structured tool interfaces. Collectively, these systems established key principles: looped reasoning, standardized tool I/O, and trajectory-based learning, yet remain geospatially naive, lacking coordinate awareness, scale handling, and domain-specific spatial verification essential for earth observation analysis.
@@
An emerging wave of EO research is now bringing these agentic principles into geospatial contexts. ThinkGeo [shabbir2025thinkgeo] frames EO question answering as tool-augmented reasoning, benchmarking ReAct-style chains and exposing persistent weaknesses in coordinate consistency, spatial grounding, and multi-step planning. Earth-Agent [feng2025earth] broadens the tool ecosystem to include spectral products and standardized interfaces but still relies on largely predefined workflows, limiting physically verifiable GIS and index-based reasoning. Geo-OLM [stamoulis2025geo] explores plan-tool-verify prompting for compact models, improving efficiency but depending primarily on prompt-level heuristics rather than learned adaptive policies. Building on this gap, we introduce an EO-native agentic stack that unifies dataset construction, supervised reasoning alignment, and evaluation under a physically grounded framework. OpenEarthAgent enables broad orchestration of GIS, index-based, optical, and SAR operations through structured instruction-tool pairs; trains on detailed reasoning traces with a dedicated evaluation set; and supports interpretable, verifiable, multi-step reasoning across diverse geospatial conditions.
3 OpenEarthAgent
3.1 Dataset Curation
@@
ReAct [yao2022react]
OpenThinkIMG [su2025openthinkimg]
ThinkGeo [shabbir2025thinkgeo]
Earth-Agent [feng2025earth]
Geo-OLM [stamoulis2025geo]
RS-Agent [xu2024rs]
RS-ChatGPT [guo2024remote]
@@
DIOR [li2020object] B-Box, Category 0.5–30 2024 Transport & Aviation Monitoring; Infrastructure Mapping
DOTA [xia2018dota] GSD, B-Box, Category 0.1–1 2021 Transport Surveillance; Infrastructure Assessment
NWPU-VHR-10 [wang2018multiscale] B-Box, Category 0.5–2 2023 Aviation; Transport; Infrastructure
UCAS-AOD [zhu2015orientation] B-Box, Category 0.5–2 2015 Aviation; Transport
AID [xia2017aid] B-Box, Category 0.2–2 2017 Urban Planning; Industrial Site Detection
iSAID [waqas2019isaid] GSD, B-Box, Seg. Map, Pix. Count 0.1–1 2019 Segmentation; Transport Monitoring
xBD [gupta2019xbd] GSD, B-Box, Category, Pix. Count 1–3.5 2019 Disaster Assessment; Change Detection
FloodNet [rahnemoonfar2021floodnet] GSD, B-Box, Category, Seg. Map, Pix. Count 0.015–0.02 2020 Flood Monitoring; Damage Estimation
Global-Dumpsite [sun2023revealing] B-Box, Category 0.3–0.8 2023 Environmental Monitoring; Waste Localization
@@
SSDD [zhang2021sar] B-Box, Category 1–15 2021 Ship Detection; Maritime Transport
SADD [zhang2022sefepnet] B-Box, Category 0.5–3 2022 Aviation; Vessel Identification
SIVED [lin2023sived] B-Box, Category 0.1–0.3 2023 Vehicle Detection; Transport Monitoring
@@
OpenStreetMap [OpenStreetMap] POIs (Points of Interest) N/A 2017- Urban Mapping; Infrastructure Analysis
@@
Google Earth Engine [GoogleEarthEngine] Index Layers 10–30 2017- Urban Planning; Change Detection; Environmental Monitoring
@@
Data Sources and Composition: The dataset integrates heterogeneous imagery and metadata from open-access EO repositories, encompassing optical, SAR, GIS, and multispectral domains. Primary sources include high-resolution benchmarks such as DOTA [xia2018dota], DIOR [li2020object], xBD [gupta2019xbd], AID [xia2017aid], NWPU-VHR-10 [wang2018multiscale], FloodNet [rahnemoonfar2021floodnet], and Global-Dumpsite [sun2023revealing], complemented by Sentinel collections and Google Earth Engine archives (see Table 2 for details). Each dataset offers distinct spatial resolutions and annotation formats (bounding boxes, segmentation masks, category labels), providing both fine-grained object understanding and large-scale semantic context. The collection spans seven thematic domains: urban planning, disaster assessment, environmental monitoring, transportation, aviation, recreation, and industrial infrastructure. Multispectral and indexed layers (e.g., NDVI, NBR, NDBI) are derived through Google Earth Engine to complement optical and SAR imagery with physically meaningful environmental indicators. Where required for spatial reasoning tasks, samples are associated with geospatial metadata such as coordinates, projection reference, and ground sampling distance (GSD), enabling verifiable geospatial analysis.
@@
(a) Data distribution across categories.
(b) Key statistics for the curated corpus used for training and evaluation of our proposed agentic pipeline.
Figure 4: Dataset overview: category distribution (left) and corpus statistics (right).
Dataset Statistics: Fig. 4 summarizes the scale, distribution, and structure of the OpenEarthAgent dataset. The training split contains 14,538 instances, and the benchmark split contains 1,169 evaluation samples. Each instance includes multimodal imagery, a natural-language query, and a structured reasoning trace with interleaved tool calls and intermediate observations. Across the training set, the corpus contains 100,656 reasoning steps (average 6.92 per query)and the benchmark split contains 7,064 reasoning steps (average 6.04 per query). Reasoning steps correspond to explicit thought-action-observation transitions, while tool calls capture the diversity of operational pathways required for solving curated agentic tasks. The breadth of modalities, thematic coverage, and depth of reasoning traces enable rigorous evaluation of grounded, interpretable, and tool-augmented geospatial reasoning.
3.2 Methodology
3.2.1 Overview:
@@
3.2.2 Unified Tool Registry:
@@
3.2.3 Reasoning Trajectories and Training Objective
@@ 4.1 Implementation Details
OpenEarthAgent is finetuned using Qwen3-4B-Instruct-2507 as the base model. Training was performed on LLM controller on four NVIDIA A100 GPUs (40 GB). The base model is loaded via Unsloth’s FastLanguageModel, enabling efficient multi-GPU distributed training. The model is trained for one epoch with a learning rate of $2\times 10^{-5}$, cosine learning rate scheduling, 0.05 warmup ratio, and a batch size of 16. The maximum sequence length is set to 4096 tokens to support long-context conversational tool interactions.